x HUBBLE SPACE TELESCOPE DAILY REPORT #1874 PERIOD COVERED: 0000Z (UTC) 05/08/97 - 0000Z (UTC) 05/09/97 Daily Status Report as of 129/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of WF/PC-2 6346 (Snapshot Survey of the = Ultraluminous IRAS Galaxy Sample-Part 1 High) The WF/PC-2 was used to make snapshot survey observations of the = ultraluminous IRAS galaxies CL-ULIR003, IR17208-00 and QDOTULIR20. These = data will be searched for multiple cores, which imply a merger has = occurred, and for starburst regions. The observations were executed as = scheduled, and no anomalies were noted. 1.2 Completed WF/PC-2 6363 (The Environments and Host Galaxies of = BL-Lac Objects) The WF/PC-2 was used to make snapshot survey observations of = BL-Lac object H2356-309 in order to study its environment and its host = galaxy. The goal is to investigate systematic trends in host galaxy = type, unusual morphological features, nearby companion galaxies, etc. = The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed WF/PC-2 6604 (Photometry of the Oldest Field Population = in the Magellanic Clouds.) The WF/PC-2 was used to make both visual and inferred photometric = observations of selected stars in the oldest field in the Large = Magellanic Cloud LMC-062630. From these data, the ages and population = densities of the cluster will be measured. The observations were = completed as scheduled, and no problems were noted. 1.4 Completed FOC/96 6618 (A Hidden Power Source in the QSO PG = 1630+377) The Faint Object Camera (f/96) was used to make observations of = quasar PG1630+337 in order to study its hidden power source. For this = hidden source to be so bright (seen only through scattering) indicates = that it may be intrinsically the most significant power source shortward = of the Lyman edge. An internal flat calibration image was also taken. = The observations were executed as scheduled, and no anomalies were noted 1.5 Completed WF/PC-2 6802 (Cosmology With the Medium Deep Survey) The WF/PC-2 was used to make observations taken in parallel with = the prime science observations as part of the Medium Deep Sky Survey. = The observations were executed as scheduled, and no problems were = reported. 1.6 Completed WF/PC-2 6428 (The Nature of the Density Cusp in the Core = of Ursa Minor) The WF/PC-2 was used to make observations of DD0199 Ursa Minoris = in order to study the cusps located in its core. Recent studies of a = central cusp in the stellar surface density profile strongly suggest the = presence of a black hole. The observations were completed as scheduled, = and no anomalies were reported. 1.7 Completed WF/PC-2 6361 (High Resolution Imaging of X-Ray Selected = AGNs) The WF/PC-2 was used to make high resolution observations of the = known, low, red-shifted, x-ray, active galactic nucleus (AGN) MS1327.4+3. = These data will allow a study of the optical properties of an unbiased = sample of AGN host galaxy. The observations were completed as scheduled, = and no problems were reported. 1.8 Completed WF/PC-2 6359 (Core Properties of the Bulges of Spiral = Galaxies) The WF/PC-2 was used to make snapshot survey observations of = spiral galaxy ESO443G80 in order to study its central bulge. T he aim is = to investigate the core properties, correlations with global galactic = properties, and the frequency of occurrence of multiple and/or = miscentered nuclei. The observations were completed as scheduled, and no = anomalies were noted. 1.9 Completed NIC/2 7154 (NICMOS Coronographic Hole Monitor) The Near Infrared Camera (NIC Two) was used to make observations = of object NGC-3602 in order to monitor the effects of coronagrapghic = hole. The purpose of this proposal is to take lamp flats with camera Two = (and cameras One and Three occasionally) to monitor the location of the = coronographic spot. The observations were completed as scheduled, and no = anomalies were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 9 Successful : 9 Scheduled Reacquisitions : 8 Successful : 8 2.2 FHST Updates: Scheduled : 21 Successful : 21 A loss of lock, described in HSTAR 6055, occurred at 129/0524Z. 2.3 Operations Notes: There were no significant problems 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science and SMOV operations. /DMH